Average height of men is normally distributed with mean height of 160 cm and standard deviation of 5 cm. If a man is randomly selected from this population, find the probability that he is between 150 cm and 162 cm. Question 12 options: 0.0228 0.3422 0.6554 0.6326 0.850
Answer #1

µ = 160

sd = 5

P(150 < X < 162)=P(150 — X- 162

                                          150 -160 162 - 1600

                                          = P(-2 < Z < 0.4)

                                          = P(Z < 0.4) - P(Z < -2)

                                          = 0.6554 - 0.0228

                                          = 0.6326 (ans)